# Robust Multi-Agent Decision-Making in Finite-Population Games

### v1

@@ IV Analysis of the KLD-RL Protocol
  • Effect of $\lambda$: Decreasing the revision rate $\lambda$ can reduce the magnitude of the noise term $\lambda(\epsilon_{i}^{N}(t)+\tilde{v}_{i}(t))$ in the closed-loop model (12c), and is therefore expected to enhance convergence to equilibrium. However, as will be discussed in Theorem 1, lowering $\lambda$ may also amplify the overall impact of noise on agents’ strategy selection. To address this, proper tuning of $\eta$ can help compensate for the increased sensitivity introduced by a smaller $\lambda$.
@@
To study how the parameters $\lambda$ and $\eta$ influence the long-term behavior of $q(t)$, we consider the closed-loop model (12) under Assumption 1. Let $(q^{\ast},x^{\ast})\in\mathbb{R}_{\geq 0}^{n}\times\mathbb{X}$ denote the unique equilibrium point of (12) in the absence of noise, where the reference distribution satisfies $\theta=x^{\ast}$. At this equilibrium, $\|q^{\ast}\|_{\infty}$ corresponds to the minimum achievable value of $\limsup_{t\to\infty}\|q(t)\|_{\infty}$, and $x^{\ast}$ corresponds to the optimal stationary distribution over the available strategies. Given that the game model is $\delta$-antipassive, [14, Theorems 5] ensures that the state of the closed-loop (12), with $\theta=x^{\ast}$, converges to a unique equilibrium $(q^{\ast},x^{\ast})$ in the absence of noise.
Throughout our analysis, we assume that the KLD-RL protocol (2) in the closed-loop model (12) satisfies $\theta=x^{\ast}$. Under this setting, the equality $C^{\eta,\theta}(p(t))-X^{N}(t)=0$ implies that agents are selecting the strategies that maximize the regularized average payoff $p^{\prime}(t)X^{N}(t)-\eta\mathcal{D}(X^{N}(t)\,\|\,x^{\ast})$. This reflects a preference for tasks with the highest remaining jobs – as captured by $p(t)=q(t)$ – while incorporating a regularization term that penalizes deviation from the optimal strategy distribution $x^{\ast}$. In addition, when the game model satisfies $\mathcal{F}(q(t),X^{N}(t))=w$, the state $X^{N}(t)$ matches the equilibrium $x^{\ast}$ for any $\eta>0$, and $\|q(t)\|_{\infty}$ attains its minimum value at $\|q^{\ast}\|_{\infty}$. Therefore, the deviation of $X^{N}(t)$ from $C^{\eta,\theta}(p(t))$ can be interpreted, in part, as a key indicator of how effectively agents are adapting the strategies toward the optimal strategy distribution $x^{\ast}$.
@@
Consider the closed-loop model (12), and recall the $\delta$-storage function $\mathcal{S}$ and $\delta$-antistorage function $\mathcal{L}$, defined for EDM (5) and the game model (7), respectively. For any finite time horizon $T>0$, the following inequality holds:
$\frac{\int_{0}^{T}\|C^{\eta,\theta}(p(t))-\widehat{X}^{N}(t)\|_{2}^{2}\,\mathrm{d}t}{\alpha_{\lambda}\!+\!\int_{0}^{T}|g_{\lambda}(p(t),\widehat{X}^{N}(t),\tilde{w}^{N}(t),\epsilon^{N}(t),\tilde{v}(t))|\mathrm{d}t}\!\leq\!(\lambda^{2}\eta)^{-1},$ (16)
@@
where $M_{\nabla_{x}\mathcal{S}}$, $M_{\mathcal{F}}$, $M_{\nabla_{q}\mathcal{L}}$, and $M_{\mathcal{V}}$ are positive constants that bound $\nabla_{x}\mathcal{S}$, $\mathcal{F}$, $\nabla_{q}\mathcal{L}$, and $\mathcal{V}^{\eta,\theta}$, respectively. [^3] Given Assumption 1 and the boundedness of $\mathbb{X}$, such constants are guaranteed to exist. Importantly, although reducing $\lambda$ lowers this upper bound, it also increases the term $(\lambda^{2}\eta)^{-1}$, thereby amplifying the effect of noise in the closed-loop model. To balance this trade-off, increasing $\eta$ can help compensate for the smaller $\lambda$. In particular, a larger $\eta$ not only reduces $(\lambda^{2}\eta)^{-1}$, but also directly decreases $\tilde{v}(t)$, due to the Lipschitz continuity of $C^{\eta,\theta}$, as shown in (17), provided that $\hat{p}(t-d)$ remains close to $p(t)$.
The remaining question is: Does strategy selection continue to improve as $\eta$ increases indefinitely? Recall that $x^{\ast}$ denotes the unique equilibrium of (12) in the absence of noise. When the population size $N$ is finite, the population state $X^{N}(t)$ cannot attain $x^{\ast}$ exactly unless $x^{\ast}\in\mathbb{X}^{N}$. Consequently, if the agents cannot reach the equilibrium state exactly, they must continuously revise their strategies in an effort to minimize $\|q(t)\|_{\infty}$.
Under Assumption 1, the function $C^{\eta,\theta}(p(t))$ approaches $x^{\ast}$ as $\eta$ becomes sufficiently large. In the analysis that follows, we show that when $\eta$ is excessively large, the evolution of $X^{N}(t)$ becomes independent of the payoff vector $p(t)$, effectively rendering (12) open-loop. As a result, (12c) becomes decoupled from (12a). To facilitate the analysis, we consider the case with no time delay in strategy revision (i.e., $d=0$).
@@
$P_{xy}^{N}$ =$=\begin{cases}x_{i}x_{j}^{\ast}&\text{if }y=x+\frac{1}{N}(e_{j}-e_{i}),~{}j\neq i\\ \sum_{i=1}^{n}x_{i}x_{i}^{\ast}&\text{if }y=x\\ 0&\text{otherwise}.\end{cases}$ (18)
@@
The proof of the proposition is presented in Appendix C. Proposition 1 implies that when $\eta$ is overly large, the population state $X^{N}(t)$ becomes independent of $p(t)$ and evolves according to $\mu^{N}$, whose variance increases as the population size $N$ decreases. As illustrated in Fig. 2, this variance can substantially influence the variability of $\|q(t)\|_{\infty}$. In conjunction with Theorem 1, the proposition implies careful tuning of $\lambda$ and $\eta$ is essential to enhance robustness to noise and convergence performance.
@@ V Conclusions
In this work, we investigated the robustness of the KLD-RL model in finite-population games, where agents select strategies based on noisy estimates of the payoff vector and revise their strategies with a time delay. Through theoretical analysis, we examined how two key parameters of the model influence the sensitivity of the strategy selection process to noise. The results provide insights into how these parameters can be tuned to enable agents to effectively learn the optimal equilibrium in task allocation games. As a direction for future research, we aim to develop data-driven approaches for optimizing the model’s parameters. In the current study, we assume that the optimal population state $x^{\ast}$ is known and that the reference distribution $\theta$ in the KLD-RL model can be set to $\theta=x^{\ast}$. To relax this assumption, we plan to explore methods for learning $x^{\ast}$ and determining a suitable choice for $\theta$.
@@
$=\lambda^{2}\nabla_{x}^{\prime}\mathcal{S}(p,x)\mathcal{V}^{\eta,\theta}(p,x)+\lambda^{2}\nabla_{x}^{\prime}\mathcal{S}(p,x)(\epsilon^{N}+\tilde{v})$
@@
$\qquad+\lambda(\lambda\nabla_{x}\mathcal{S}(p,x)+\mathcal{F}(q,x)-w)^{\prime}(\epsilon^{N}+\tilde{v})$
@@
Following the analysis in [18], define the log-sum-exp function $\mathrm{lse}^{\eta,\theta}:\mathbb{R}^{n}\to\mathbb{R}$ associated with (2) as:
@@
where $\mathrm{diag}(C^{\eta,\theta}(p))$ denotes the diagonal matrix whose diagonal entries are given by the components of $C^{\eta,\theta}(p)$. Consequently, by following similar steps as in the proof of [18, Proposition 4], it holds for any $p,v\in\mathbb{R}^{n}$ that
@@
Using [18, Lemma 5], we conclude the function $C^{\eta,\theta}$ is Lipschitz continuous:

### v2

@@ IV Analysis of the KLD-RL Protocol
  • Effect of $\lambda$: When $\epsilon_{i}^{N}(t)$ and $\tilde{v}_{i}(t)$ are fixed, the magnitude of the noise term $\lambda(\epsilon_{i}^{N}(t)+\tilde{v}_{i}(t))$ in the closed-loop model (12c) scales linearly with $\lambda$. Hence, decreasing $\lambda$ would reduce the noise level and promote convergence to equilibrium. However, as shown in Theorem 1, a smaller $\lambda$ may also amplify the overall impact of noise on agents’ strategy selection. This effect can be alleviated through proper tuning of $\eta$, which helps counterbalance the increased sensitivity introduced by a reduced $\lambda$.
@@
To study how the parameters $\lambda$ and $\eta$ affect the long-term behavior of $q(t)$, we consider the closed-loop model (12) under Assumption 1. Let $(q^{\ast},x^{\ast})$ denote the unique equilibrium point of (12) in the absence of noise, where the reference distribution satisfies $\theta=x^{\ast}$. At this equilibrium, $\|q^{\ast}\|_{\infty}$ corresponds to the minimum achievable value of $\limsup_{t\to\infty}\|q(t)\|_{\infty}$, and $x^{\ast}$ corresponds to the optimal stationary distribution over the available strategies. Given that the game model is $\delta$-antipassive, [9029756, Theorems 5] ensures that the state of the closed-loop (12), with $\theta=x^{\ast}$, converges to $(q^{\ast},x^{\ast})$ in the absence of noise.
Throughout our analysis, we assume that the KLD-RL protocol (2) in the closed-loop model (12) satisfies $\theta=x^{\ast}$. Under this setting, the equality $C^{\eta,\theta}(p(t))-X^{N}(t)=0$ implies that agents are selecting the strategies that maximize the regularized average payoff $p^{\prime}(t)X^{N}(t)-\eta\mathcal{D}(X^{N}(t)\,\|\,x^{\ast})$. This reflects a preference for tasks with the highest remaining jobs—as captured by p(t)=q(t)p(t)=q(t)—while incorporating a regularization term that penalizes deviation from the optimal strategy distribution $x^{\ast}$. In addition, when the game model satisfies $\mathcal{F}(q(t),X^{N}(t))=w$, the state $X^{N}(t)$ matches the equilibrium $x^{\ast}$ for any $\eta>0$, and $\|q(t)\|_{\infty}$ attains its minimum value at $\|q^{\ast}\|_{\infty}$. Therefore, the deviation of $X^{N}(t)$ from $C^{\eta,\theta}(p(t))$ can be interpreted, in part, as a key indicator of how effectively agents are adapting the strategies toward the optimal strategy distribution $x^{\ast}$.
@@
Consider the closed-loop model (12), and recall the $\delta$-storage function $\mathcal{S}$ and $\delta$-antistorage function $\mathcal{L}$, defined for EDM (5) and the game model (7), respectively. For any time horizon $T>0$, the following inequality holds:
$\textstyle\int_{0}^{T}\|C^{\eta,\theta}(p(t))-\widehat{X}^{N}(t)\|_{2}^{2}\,\mathrm{d}t$
$\textstyle\leq\!\frac{1}{\lambda^{2}\eta}\!\left(\alpha_{\lambda}\!+\!\int_{0}^{T}\!|g_{\lambda}(p(t),\widehat{X}^{N}(t),\tilde{w}^{N}(t),\epsilon^{N}(t),\tilde{v}(t))|\mathrm{d}t\right),$ (16)
@@
where $_{\nabla_{x}\mathcal{S}}, MM_{\mathcal{F}}$, $M_{\nabla_{q}\mathcal{L}}$, and $M_{\mathcal{V}}$ are positive constants that bound $\nabla_{x}\mathcal{S}$, $\mathcal{F}$, $\nabla_{q}\mathcal{L}$, and $\mathcal{V}^{\eta,\theta}$, respectively. [^4] Given Assumption 1 and the boundedness of $\mathbb{X}$, such constants are guaranteed to exist. Notably, although reducing $\lambda$ lowers this upper bound, it also increases the term $(\lambda^{2}\eta)^{-1}$, thereby amplifying the effect of noise in the closed-loop model. To balance this trade-off, increasing $\eta$ can help compensate for the smaller $\lambda$. In particular, a larger $\eta$ not only reduces $(\lambda^{2}\eta)^{-1}$, but also directly decreases $\tilde{v}(t)$, due to the Lipschitz continuity of $C^{\eta,\theta}$, as shown in (17), provided that $\hat{p}(t-d)$ remains close to $p(t)$.
The remaining question is: Can agents increase $\eta$ arbitrarily to compensate for a smaller $\lambda$, thereby further mitigating the influence of noise on strategy selection?
Recall that $x^{\ast}$ denotes the unique equilibrium of (12) in the absence of noise. Under Assumption 1, the function $C^{\eta,\theta}(p(t))$ converges to $x^{\ast}$ as $\eta$ becomes sufficiently large. In the analysis that follows, we show that when $\eta$ is excessively large; hence, we approximate $C^{\eta,\theta}(p(t))$ with $x^{\ast}$, the evolution of $X^{N}(t)$ becomes independent of the payoff vector $p(t)$, effectively rendering (12) open-loop. As a result, (12c) becomes decoupled from (12a). To facilitate the analysis, we consider the case with no time delay in strategy revision (i.e., $d=0$).
@@
$P_{xy}^{N}$ $=\begin{cases}x_{i}x_{j}^{\ast}&\text{if }y=x+\frac{1}{N}(e_{j}-e_{i}),~j\neq i\\ \sum_{i=1}^{n}x_{i}x_{i}^{\ast}&\text{if }y=x\\ 0&\text{otherwise}.\end{cases}$ (18)
@@
The proof of the proposition is provided in Appendix C. Proposition 1 indicates that when $\eta$ is excessively large, the population state $X^{N}(t)$ becomes independent of $p(t)$ and evolves according to $\mu^{N}$, whose variance increases as the population size $N$ decreases. As shown in Fig. 2, in this regime the agents cannot adaptively switch strategies based on $p(t)$, leading to overshooting in the early stages. Furthermore, the variance of $X^{N}(t)$ may substantially affect the variability of $\|q(t)\|$ in later periods. Taken together with Theorem 1, the proposition highlights the importance of carefully tuning $\lambda$ and $\eta$ to improve both noise robustness and convergence performance.
@@ V Conclusions
We investigated the robustness of the KLD-RL model in finite-population games, where agents choose strategies based on noisy payoff estimates and update their strategies with a time delay. We analyzed how two key parameters affect the sensitivity of strategy selection to noise. Our results highlight how tuning these parameters enables agents to effectively learn the optimal equilibrium in task allocation games. For future work, we aim to develop data-driven methods for optimizing the model’s parameters. In this study, we assumed that the optimal population state $x^{\ast}$ is known and set the model’s reference distribution to $\theta=x^{\ast}$. To relax this assumption, we plan to explore approaches for learning $x^{\ast}$ directly and identifying an appropriate choice for $\theta$.
@@
[…]
@@
[…]
@@
Following the analysis in [gao2018propertiessoftmaxfunctionapplication], define the log-sum-exp function $\mathrm{lse}^{\eta,\theta}:\mathbb{R}^{n}\to\mathbb{R}$ associated with (2) as:
@@
where $\mathrm{diag}(C^{\eta,\theta}(p))$ denotes the diagonal matrix whose diagonal entries are given by the components of $C^{\eta,\theta}(p)$. Consequently, by following similar steps as in the proof of [gao2018propertiessoftmaxfunctionapplication, Proposition 4], it holds for any $p,v\in\mathbb{R}^{n}$ that
@@
Using [gao2018propertiessoftmaxfunctionapplication, Lemma 5], we conclude the function $C^{\eta,\theta}$ is Lipschitz continuous: